RC meeting and Run Plan, Wed Oct. 10

From HallCWiki
Jump to: navigation, search
  • Status
    • Data taken so far [1]
    • Beam quality [2]
      • Beam issues, no data until 21:40pm
      • Lost ~90minutes due to beam dump water flow alert (see below).
  • Requiring attention
    • Around midnight beam dump water flow alert (flow too high) [3]
      • Ultimately resolved by masking [4]
      • BUT... around 10:30 downtime created for the same alert [5]
      • Again resolved by masking ...after 49 minutes
      • Need to make sure this gets communicated to the next crew if needed
    • RadCon wants to put a small detector in the SHMS hut during the next opportunity
      • Needs few hours, communicate with Brad when next opportunity arrises
  • Plans - Physics
    • Take data at Q2=3. W=3.14 x=0.25 theta_Shms=9.42 to reach 66% of desired stat. goal., then take Al.
      • Projected to take until tomorrow 4pm
    • Change SHMS angle to 6.48 deg, take Al and LH2 data
      • Coordinate with Holly in case we trigger rotation alarm.
      • Coordinate with Jay to see if we still need new limits for the ion chamber diffusors at this angle [6]
      • Projected to put us sometime in the weekend
    • Start preparation for 2-shift luminosity scan
      • Will need stable beam for this, and experienced shift
      • Communicate with Jay Benesch to find best time, might have to be opportunistic